There are n persons and m monkeys (m>n) no of ways in which each person may become the owner of one monkey is

Grade:11

1 Answers

Rishabh Ranjan
6 years ago
The answer to this question is m-1Cn-1
By using multinomial theorem, we can find this out
a1>=1,a2>=1,a3>=1,...,an>=1
a1+a2+a3+...+an=m – n
Therefore, the no. of ways in which each person may become the owner of one monkey is = m-n+n-1Cn-1 = m-1Cn-1